# On the Reachability of 3–Dimensional Paths with a Prescribed Curvature Bound

### v2

@@ Abstract
This paper presents the reachability analysis of curves in $\mathbb{R}^{3}$ with a prescribed curvature bound. Based on Pontryagin–Maximum–Principle, internal connection between time–optimal control problems and reachability is discussed. This finding allows for leveraging the existing knowledge on the structure of solutions to minimum–time problems, or the Markov–Dubins problem, to reachability considerations. Based on this development, two types of reachability are discussed. First, we prove that any boundary point of the reachability set, with the directional component taken into account as well as geometric coordinates, can be reached via curves of H, CSC, CCC, or their respective subsegments, where H denotes a helicoidal arc, C a circular arc with maximum curvature, and S a straight segment. This provides an implicit description of the reachability set. Second, we show that any boundary point of the reachability set when directional component is not considered—the position reachability set—can be reached by planar curves of CS, CC, or their subsegments. In this case, we provide complete description of the reachability set by proving that it is a solid of revolution of its two–dimensional counterpart, the Dubins car. These findings extend the developments presented in literature regarding planar curves, or Dubins car dynamics, into spatial curves in $\mathbb{R}^{3}$. These advancements in understanding the reachability of curvature bounded paths in $\mathbb{R}^{3}$ hold significant practical implications, particularly in the contexts of path planning and time optimal guidance.
@@ 1 Introduction
The analysis of reachability, or attainability, of dynamical systems is crucial in a wide range of applications. This includes solving optimal control problems (OCPs) (McGregor et al. (1999)) and collision–avoidance path planning strategies (Althoff (2010)). With such background, various studies have explored methods for constructing the boundary of the reachability sets of curves (paths) in two–dimensional (2D) plane with prescribed curvature bound (Melzak (1961)Wong and Korsak (1974)Cockayne and Hall (1975)Patsko et al. (2003)Patsko and Fedotov (2022)). The emphasis on the curvature–bounded paths is due to the fact that the dynamics of planar curves with a prescribed curvature bound, also known as the Dubins car (Dubins (1957)), represents numerous practical applications. These include scenarios involving fixed–wing unmanned aerial vehicles (UAVs) (Lin and Saripalli (2014)), missile guidance problems (Zheng et al. (2021)), and autonomous underwater vehicles (AUVs) (Scibilia et al. (2012)).
@@
Many studies on the reachability of Dubins cars stem from the foundational work by L. E. Dubins (Dubins (1957)). In this work, Dubins addressed the 2D case of the problem first introduced by A. A. Markov, which questions the structure of the shortest paths with prescribed curvature bounds, along with specified initial and terminal locations and directions. Also known as Markov–Dubins problem in general dimensions, the 2D case was first solved in Dubins (1957). Dubins proved that the minimizers are curves of CSC, CCC, or their subsegments, where C denotes a circular arc with maximum curvature and S a straight segment. (Throughout this paper, we will refer to the collection of all CSC curves as class CSC. This notation will be applied strictly in a sense that, for instance, any CSC curve with the S component having zero length will be referred to as CC, not CSC. Subsegments of, for instance, class CSC, refer to classes C, S, CS, SC, and CC.) While the central arguments in Dubins (1957) were based on geometric approaches, subsequent works have reformulated the problem as an OCP and applied the Pontryagin Maximum Principle (PMP) to reproduce the same results (Johnson (1974)Boissonnat et al. (1992)).
@@
These works, which solved the minimum time problem in terms of optimal control, facilitated subsequent studies on the reachability of Dubins car. The question was as follows (Cockayne and Hall (1975)): “A particle moves in $\mathbb{R}^{2}$ with constant speed and subject to an upper bound on the curvature of its path. At each time $t$, what is the set of all possible positions for the particle?” The small–time case was addressed in Melzak (1961), followed by Cockayne and Hall (1975) for general $t$. By applying geometric arguments, the authors of Cockayne and Hall (1975) first introduced an implicit description of the reachability set by proving that every boundary point of the reachability set can be reached by curves in classes CS, CC, and their subsegments. Hence, collecting the endpoints of all such curves provides a small set containing the boundary. Then, an additional condition that the reachability set varies continuously in time facilitated complete and analytical description of the boundary of the reachability set. In this problem, the target of reachability analysis was only the geometric coordinates, which we refer to as position reachability set.
@@
Reachability analysis of Dubins car when the directional components are also considered—which we refer to as reachability set without any modifier—was first addressed in Patsko et al. (2003). The authors used PMP to conclude that every boundary point of the reachability set can be reached by curves in classes CSC, CCC, or their subsegments. This implicit description does not solely provide complete description of the boundary of the reachability set for general $t$. However, it was sufficient for construction of the reachability set for small–time. Subsequently, a follow–up study in Patsko and Fedotov (2022) was conducted to provide additional conditions to facilitate complete and analytical description of the reachability set for all $t$.
The prior developments on the reachability of the Dubins car (Cockayne and Hall (1975)Patsko et al. (2003)) have enabled numerous subsequent advancements in multiple fields. To illustrate, the analytical description of the position reachability set in Cockayne and Hall (1975) facilitated the development of a path planning method under the presence of dynamic obstacles studied in Wu and How (2012) for guaranteed collision avoidance. It also enabled the computation of the shortest paths to a given straight line segments in Vendittelli et al. (1999) for obstacle avoidance. Studies on time–optimal interception in Buzikov and Galyaev (2021) and Buzikov (2024) were grounded on this development as well. The implicit description of the reachability set of Dubins car in Patsko et al. (2003), although it was not a complete description, also served as a foundational groundwork for multiple studies. These include Buzikov and Galyaev (2022) on time–optimal interceptions with constrained interception angles, the path elongation method developed in Chen et al. (2023), and the complete and analytical description of the reachability set in Patsko and Fedotov (2022).
@@
Acknowledging this impact of reachability analysis of curvature bounded paths, we extend the existing results on Dubins car into 3D curves with prescribed curvature bound. Since names such as Dubins airplane or 3D Dubins car have already been taken in the literature to represent different dynamics—an extension of Dubins car into 3D by including an independent altitude component (Chitsaz and LaValle (2007))—we will refer to our dynamics as 3D curves with prescribed curvature bound. Unlike ‘Dubins airplane’, our dynamics maintains the curvature bound property, ensuring that the curvature of the trajectory remains tractable.
As reachability analysis of Dubins car in literature utilized developments from its minimum time problem, we similarly leverage results of the minimum time problem for the 3D case. The Markov–Dubins problem in 3D was addressed in Sussmann (1995) using PMP, and Hota and Ghose (2010) provides numerical methods to compute the minimizer when initial and terminal locations are far enough. Our work is grounded on formulation of PMP on manifolds, and the developments made in Sussmann (1995) will be leveraged extensively.
@@
The contributions of this paper are summarized as follows: First, we reveal the internal connection between reachability and time optimality through application of PMP. This is facilitated by introducing set of OCPs that have the same PMP conditions. This allows to apply useful conclusions made in Sussmann (1995) on the minimum time problem to our reachability considerations. Second, we present implicit description of the reachability set of 3D curves with prescribed curvature bound. (i.e., when directional components are also considered.) We prove that the boundary points of the reachability set can be reached via CCC, CSC, their subsegments, or H, where H denotes certain class of helicoidal arcs. This extends the results made in Patsko et al. (2003) into dimension 3. As with Patsko et al. (2003), our development provides an implicit description. Given the significant benefits drawn from implicit descriptions in previously cited works, we believe this will make a valuable contribution to the field. Last, we present complete and analytical description of the reachability set of 3D curves with prescribed curvature bound. We prove that the position reachability set in this case can be simply generated by rotating the position reachability set of Dubins car in Cockayne and Hall (1975). This extends the results made in Cockayne and Hall (1975) into dimension 3. These developments in this paper are poised to inspire numerous future studies by extending the existing literature from 2D to 3D scenarios.
The remaining sections of this paper are organized as follows. Section 2 introduces definitions and preliminary results in literature. OCP formulation of the Markov Dubins problem in dimensions 2 and 3 are outlined alongside. Section 3 first proves the internal relationship between time optimal problems and reachability set. Then, the developments in Cockayne and Hall (1975) and Patsko et al. (2003) are extended into dimension 3. Section 4 presents the concluding remarks of this paper.
@@ 2.1 Definitions and preliminary theorems
  • For each $\boldsymbol{u}\in\mathcal{F}$, there exists a response on $[0,T]$.
@@
  • For each $\boldsymbol{x}$, the set $V(\boldsymbol{x})=\{f(\boldsymbol{x},\boldsymbol{u}):\boldsymbol{u}\in\Omega\}$ is convex.
In this context, uniform bound means that the choice of the upper bound $m$ is independent of the choice of controller $\boldsymbol{u}\in\mathcal{F}$ and time $t$. Compactness of $V(\boldsymbol{x})$ follows trivially from compactness of $\Omega$ and continuity of $f$. Assumption 1 is sufficient for the discussions introduced in this paper, but not all of them are necessary everywhere. Indeed, some can be relaxed, although we will not attempt fine–tuning beyond the purpose of applications to 3D curves with prescribed curvature bounds.
For $t\in[0,T]$, the Reachability Set $\mathcal{G}(t)$ is defined as follows.
@@
Assumption 1 and the well–known Filippov’s Theorem assert that each $\mathcal{G}(t)$ is compact. Moreover, under Assumption 1, it was proved in Lee and Markus (1967) that the reachability set $\mathcal{G}(t)$ varies continuously in time on $[0,T]$, where continuity is with respect to Hausdorff metric. It was further proved in Lee and Markus (1967) the following theorem, where boundary of a set $S$ is denoted as $bd(S)$.
@@
The ‘almost everywhere’ condition is abbreviated as a.e. $t$.
A slightly stronger statement is available if all responses lie on a smooth manifold $M$. (This inherently subsumes the $\mathbb{R}^{n_{\boldsymbol{x}}}$ case.) In this case, since each $f(\boldsymbol{x}^{*},\boldsymbol{u}^{*})$ is a tangent vector in $T_{\boldsymbol{x}^{*}}M$, the adjoint response $\boldsymbol{p}^{*}$—which we take inner product with $f(\boldsymbol{x}^{*},\boldsymbol{u}^{*})$—is viewed as a covector on $M$ rather than a vector in the ambient space $\mathbb{R}^{n_{\boldsymbol{x}}}$. Consequently, the nontriviality condition of $\boldsymbol{p}^{*}$ is applied by interpreting $\boldsymbol{p}^{*}$ as a covector on $M$, instead of merely being a nonzero vector in $\mathbb{R}^{n_{\boldsymbol{x}}}$. In other words, $\boldsymbol{p}^{*}$ has to be considered as a nonvanishing linear functional $\boldsymbol{z}\mapsto\langle\boldsymbol{p}^{*},\boldsymbol{z}\rangle$ from $T_{\boldsymbol{x}^{*}}M$ to $\mathbb{R}$. Benefit of such distinction between the nontriviality conditions will be apparent in the subsequent sections. A more detailed description regarding the formulation of maximum principle under manifold setting can be found in Sussmann (1998). Under this consideration, the following theorem directly follows from Theorem 8.2 in Sussmann (1998).
@@
Definition of solution of an ordinary differential equation (ODE) is Carathéodory sense as usual. Since $\mathcal{G}(t_{f})$ is compact under Assumption 1, all of its boundary points are contained in itself and therefore reachable at time $t_{f}$. Consequently, the following corollary holds for each $t_{f}$.
@@
Additional information contained in Theorem 3 but not in Corollary 6 is Theorem 3 states that for ‘every’ trajectory reaching the boundary, there exists a nontrivial solution of Eqs. (3) and (4). Hence, the points in $E^{*}(t_{f})$ that can be reached by an alternative trajectory that only allows trivial adjoint response of Eqs. (3) and (4) is an interior point of $\mathcal{G}(t_{f})$. This idea was leveraged in the reachability analysis of Dubins car in Patsko et al. (2003) (e.g., Lemma 2).
Another information, which in this case not contained in Theorem 3, is that $\mathcal{G}(t)$ varies continuously in time. This facilitated complete description of the position reachability set of Dubins car for all $t$ (Cockayne and Hall (1975)). Throughout our reachability analysis in Section 3, we will utilize these two additional information extensively. Before presenting the main results, we introduce the Markov–Dubins problem in dimensions 2 and 3, as well as the reachability analysis of the Dubins car in the literature, for the completeness of paper.
2.2 Markov–Dubins problem in dimensions 2 and 3
Formulation of the Markov–Dubins problem is as follows: given a constant $\kappa_{M}>0$, two points $\boldsymbol{x}_{0}$, $\boldsymbol{x}_{1}\in\mathbb{R}^{n}$, and unit vectors $\boldsymbol{y}_{0}$, $\boldsymbol{y}_{1}\in\mathbb{R}^{n}$, find a curve $\gamma:[0,L]\mapsto\mathbb{R}^{n}$ of shortest length (i.e., minimum $L$) such that
  1. 1.
    $\gamma(s)$ is parametrized by arc length,
  2. 2.
    $\gamma(0)=\boldsymbol{x}_{0}$, $\gamma^{\prime}(0)=\boldsymbol{y}_{0}$, $\gamma(L)=\boldsymbol{x}_{1}$, and $\gamma^{\prime}(L)=\boldsymbol{y}_{1}$,
  3. 3.
    $\gamma^{\prime}$ is absolutely continuous and satisfies the curvature bound: $|\gamma^{\prime\prime}(s)|\leq\kappa_{M}$ for almost all $s\in[0,L]$.
The case of $n=2$, also known as Dubins car, was first studied in Dubins (1957) by geometric means proving that the minimizers are of CSC, CCC, and their subsegments. After the foundational study of Dubins, Markov–Dubins problem of $n=2$ was formulated as a minimum time optimal control problem and treated by means of PMP in Johnson (1974) and Boissonnat et al. (1992) to reproduce the same results. The problem setup was as follows, taking $\kappa_{M}=1$ for simplicity.
@@ 2.2 Markov–Dubins problem in dimensions 2 and 3
where the state variable is defined as $\boldsymbol{\chi}=[x^{1},\;x^{2},\;\theta]^{T}\in\mathbb{R}^{3}$ and control variable as $u\in\mathbb{R}$. $x^{1}$ and $x^{2}$ represent the geometric coordinates and $\theta$ represents the heading angle. Admissible controls are all measurable functions $t\mapsto u(t)\in[-1,1]$.

### v3

# On the Reachability of 3-Dimensional Paths with a Prescribed Curvature Bound

@@ Abstract
This paper presents the reachability analysis of curves in $\mathbb{R}^{3}$ with a prescribed curvature bound. Based on Pontryagin Maximum Principle, we leverage the existing knowledge on the structure of solutions to minimum-time problems, or Markov-Dubins problem, to reachability considerations. Based on this development, two types of reachability are discussed. First, we prove that any boundary point of the reachability set, with the directional component taken into account as well as geometric coordinates, can be reached via curves of H, CSC, CCC, or their respective subsegments, where H denotes a helicoidal arc, C a circular arc with maximum curvature, and S a straight segment. Second, we show that the reachability set when directional component is not considered—the position reachability set—is simply a solid of revolution of its two-dimensional counterpart, the Dubins car. These findings extend the developments presented in literature on Dubins car into spatial curves in $\mathbb{R}^{3}$.
@@ 1 Introduction
Analysis of reachability of dynamical systems is crucial in a wide range of applications (Althoff (2010)). Accordingly, various studies have explored methods for constructing the reachability set of curves in two-dimensional (2D) plane with prescribed curvature bound (Melzak (1961)Wong and Korsak (1974)Cockayne and Hall (1975)Patsko et al. (2003)Patsko and Fedotov (2022)). The emphasis on the curvature-bounded paths is because the dynamics of planar curves with a prescribed curvature bound, or Dubins car (Dubins (1957)), represents numerous practical platforms.
@@
In Dubins (1957), the author addressed the 2D case of the problem first introduced by A. A. Markov, also known as Markov-Dubins problem, which questions the structure of the shortest paths in $\mathbb{R}^{n}$ with prescribed curvature bounds along with specified initial and terminal locations and directions. Dubins proved that the minimizers are curves of CSC, CCC, or their subsegments in 2D, where C denotes a circular arc with maximum curvature and S a straight segment. (Throughout this paper, we will refer to the collection of all CSC curves as class CSC. This notation will be applied strictly in a sense that, for instance, any CSC curve with the S component having zero length will be referred to as CC, not CSC. Subsegments of, for instance, class CSC, refer to classes C, S, CS, SC, and CC.) While the central arguments in Dubins (1957) were based on geometric approaches, subsequent works have reformulated the problem as an OCP and applied the Pontryagin Maximum Principle (PMP) to reproduce the same results (Pecsvaradi (1972)Johnson (1974)Boissonnat et al. (1992)).
@@
These works facilitated subsequent studies on reachability of Dubins car. The question was as follows (Cockayne and Hall (1975)): “A particle moves in $\mathbb{R}^{2}$ with constant speed and subject to an upper bound on the curvature of its path. At each time $t$, what is the set of all possible positions for the particle?” The small-time case was addressed in Melzak (1961), followed by Cockayne and Hall (1975) for general $t$. By applying geometric arguments, Cockayne and Hall (1975) first introduced an implicit description of the reachability that every boundary point can be reached by curves in classes CS, CC, and their subsegments. Hence, collecting the endpoints of all such curves provides a small set containing the boundary. Then, an additional condition that the reachability set varies continuously in time facilitated complete and analytical description of the boundary of the reachability set. In this problem, the target of reachability analysis was only the geometric coordinates, which we refer to as position reachability set.
@@
Reachability analysis of Dubins car with directional components considered was first addressed in Patsko et al. (2003). PMP was used to conclude that every boundary point of the reachability set can be reached by curves in classes CSC, CCC, or their subsegments. This implicit description does not solely provide complete description of the boundary of the reachability set for general $t$, but was sufficient for construction for small-time. Subsequently, a follow-up study in Patsko and Fedotov (2022) provided additional conditions to facilitate complete and analytical description of the reachability set for all $t$.
@@
Prior developments on the reachability of the Dubins car (Cockayne and Hall (1975)Patsko et al. (2003)) enabled numerous subsequent advancements (Buzikov and Galyaev (2021)Buzikov (2024)Buzikov and Galyaev (2022)Chen et al. (2023)Patsko and Fedotov (2022)). Motivated by these works, we extend the existing results on Dubins car into 3D curves with prescribed curvature bound. Since names such as Dubins airplane or 3D Dubins car have already been taken in the literature to represent different dynamics—extension of Dubins car into 3D by including an independent altitude component (Chitsaz and LaValle (2007))—we will refer to our dynamics as 3D curves with prescribed curvature bound. Unlike ‘Dubins airplane’, our dynamics maintains the curvature bound property. As reachability analysis of Dubins car in literature utilized the developments on its minimum time problem, we similarly leverage the results of minimum time problem for the 3D case studied in Sussmann (1995).
@@
Contributions of this paper are as follows: First, we reveal the internal connection between reachability and time optimality which allows to apply the conclusions in Sussmann (1995) on the minimum time problem to reachability considerations. Second, we prove that the boundary points of the reachability set (i.e., when directional components are also considered.) can be reached via CCC, CSC, their subsegments, or H, where H denotes certain class of helicoidal arcs. This extends the results in Patsko et al. (2003) into dimension 3. Last, we present complete and analytical description of the reachability set of 3D curves with prescribed curvature bound. We prove that the position reachability set in this case can be simply generated by rotating the position reachability set of Dubins car in Cockayne and Hall (1975). This extends the results made in Cockayne and Hall (1975) into dimension 3.
@@ 2.1 Definitions and preliminary theorems
  • For each $\boldsymbol{u}\in\mathcal{F}$, there exists a response on $[0,T]$,
@@
  • $V(\boldsymbol{x})=\{f(\boldsymbol{x},\boldsymbol{u}):\boldsymbol{u}\in\Omega\}$ is convex for each $\boldsymbol{x}$.
In this context, uniform bound means that the choice of the upper bound $m$ is independent of the choice of controller $\boldsymbol{u}\in\mathcal{F}$ and time $t$. Compactness of $V(\boldsymbol{x})$ follows trivially from compactness of $\Omega$ and continuity of $f$. Assumption 1 is sufficient for our discussions though some can be relaxed beyond the purpose of applications to 3D curves with prescribed curvature bounds. For $t\in[0,T]$, the Reachability Set $\mathcal{G}(t)$ is defined as follows.
@@
Assumption 1 and the well-known Filippov’s Theorem assert that each $\mathcal{G}(t)$ is compact. Moreover, under Assumption 1, it was proved in Lee and Markus (1967) that $\mathcal{G}(t)$ varies continuously in time on $[0,T]$, where continuity is with respect to Hausdorff metric. It was further proved in Lee and Markus (1967) the following theorem, where boundary of a set $S$ is denoted as $bd(S)$.
@@
The ‘almost everywhere’ condition is abbreviated as a.e. $t$. A slightly stronger statement is available if all responses lie on a smooth manifold $M$. In this case, since each $f(\boldsymbol{x}^{*},\boldsymbol{u}^{*})$ is a tangent vector in $T_{\boldsymbol{x}^{*}}M$, the adjoint response $\boldsymbol{p}^{*}$ is viewed as a covector on $M$ rather than a vector in the ambient space $\mathbb{R}^{n_{\boldsymbol{x}}}$. Consequently, the nontriviality condition of $\boldsymbol{p}^{*}$ is applied by interpreting $\boldsymbol{p}^{*}$ as a covector on $M$, instead of merely being a nonzero vector in $\mathbb{R}^{n_{\boldsymbol{x}}}$. In other words, $\boldsymbol{p}^{*}$ has to be considered as a nonvanishing linear functional $\boldsymbol{z}\mapsto\langle\boldsymbol{p}^{*},\boldsymbol{z}\rangle$ from $T_{\boldsymbol{x}^{*}}M$ to $\mathbb{R}$. Benefit of such distinction between the nontriviality conditions will be apparent in the subsequent sections. A more detailed description regarding the formulation of maximum principle under manifold setting can be found in Sussmann (1998). Under this consideration, the following theorem directly follows from Theorem 8.2 in Sussmann (1998).
@@
Since $\mathcal{G}(t_{f})$ is compact under Assumption 1, all of its boundary points are contained in itself and therefore reachable at time $t_{f}$. Consequently, the following corollary holds for each $t_{f}$.
@@
Additional information contained in Theorem 3 but not in Corollary 6 is Theorem 3 states that for ‘every’ trajectory reaching the boundary, there exists a nontrivial solution of Eqs. (3) and (4). Hence, the points in $E^{*}(t_{f})$ that can be reached by an alternative controller that is not maximal (i.e., violates the condition in Eq. (4)) lie interior to $\mathcal{G}(t_{f})$. This was leveraged in the reachability analysis of Dubins car in Patsko et al. (2003) (e.g., Lemma 2).
Another information, which in this case not contained in Theorem 3, is that $\mathcal{G}(t)$ varies continuously in time. This facilitated complete description of the position reachability set of Dubins car for all $t$ (Cockayne and Hall (1975)). Throughout our reachability analysis in Section 3, we will utilize these two additional information extensively.
2.2 Markov-Dubins problem in dimensions 2 and 3
Formulation of Markov-Dubins problem is as follows: given a constant $\kappa_{M}>0$, two points $\boldsymbol{x}_{0}$, $\boldsymbol{x}_{1}\in\mathbb{R}^{n}$, and unit vectors $\boldsymbol{y}_{0}$, $\boldsymbol{y}_{1}\in\mathbb{R}^{n}$, find a curve $\gamma:[0,L]\mapsto\mathbb{R}^{n}$ of shortest length (i.e., minimum $L$) such that (1) $\gamma(s)$ is parametrized by arc length, (2) $\gamma(0)=\boldsymbol{x}_{0}$, $\gamma^{\prime}(0)=\boldsymbol{y}_{0}$, $\gamma(L)=\boldsymbol{x}_{1}$, and $\gamma^{\prime}(L)=\boldsymbol{y}_{1}$, and (3) $\gamma^{\prime}$ is absolutely continuous and satisfies the curvature bound: $|\gamma^{\prime\prime}(s)|\leq\kappa_{M}$ for almost all $s\in[0,L]$.
The case of $n=2$, also known as Dubins car, was first studied in Dubins (1957) by geometric means proving that the minimizers are of CSC, CCC, and their subsegments. After Dubins (1957), the problem was formulated as a minimum time optimal control problem and treated by means of PMP in Pecsvaradi (1972)Johnson (1974), and Boissonnat et al. (1992) to reproduce the same results. The problem setup was as follows, taking $\kappa_{M}=1$ for simplicity.
@@ 2.2 Markov-Dubins problem in dimensions 2 and 3
where the state variable is $\boldsymbol{\chi}=[x^{1},\;x^{2},\;\theta]^{T}\in\mathbb{R}^{3}$ and control variable is $u\in\mathbb{R}$. $x^{1}$ and $x^{2}$ represent the geometric coordinates and $\theta$ represents the heading angle. Admissible controls are all measurable functions $t\mapsto u(t)\in[-1,1]$.